OpenStudy (shreenithaa):

Help!!! Medal!!! Sure Print: Ticket Printing 25 tickets for $2 Best Print Tickets printed $10 setting up plus $1 for 25 tickets Using C for the cot of the printing and T for the number of tickets. Write the formula. Sure Print: C= 2T/25 Best Print: C=1T/25 +10 Draw a graph (I already did that) Use algebra to find C and T. 2T/25=1T/25 +10 T=25 Which company would you choose? Why? (I need the why to be clearly explained you can use the questions above to answer the why. The why is the most important.

OpenStudy (shreenithaa):

Sure Print: Ticket Printing 25 tickets for $2 Best Print Tickets printed $10 setting up plus $1 for 25 tickets Using C for the cot of the printing and T for the number of tickets. Write the formula. Sure Print: C= 2T/25 Best Print: C=1T/25 +10 Draw a graph (I already did that) Use algebra to find C and T. 2T/25=1T/25 +10 T=25 Which company would you choose? Why? (I need the why to be clearly explained you can use the questions above to answer the why. The why is the most important.

OpenStudy (shreenithaa):

does anyone get this?